Whispering Pines Inc. is all-equity-financed. The expected rate of return on the company’s shares is 11.45%.

a. What is the opportunity cost of capital for an average-risk Whispering Pines investment? (Enter your answer as a percent rounded to 2 decimal places.)


b. Suppose the company issues debt, repurchases shares, and moves to a 28% debt-to-value ratio (D/V = 0.28). What will be the company’s weighted-average cost of capital at the new capital structure? The borrowing rate is 7.15% and the tax rate is 21%. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

Answer a.

Opportunity cost of capital of an all-equity financed company is equal to the unlevered cost of equity.

Opportunity cost of capital = Unlevered cost of equity
Opportunity cost of capital = 11.45%

Answer b.

Weight of debt = 0.28
Weight of equity = 0.72

Debt-equity ratio = Weight of debt / Weight of equity
Debt-equity ratio = 0.28 / 0.72
Debt-equity ratio = 0.3889

Levered cost of equity = Unlevered cost of equity + (Unlevered cost of equity - Cost of debt) * Debt-equity ratio
Levered cost of equity = 0.1145 + (0.1145 - 0.0715) * 0.3889
Levered cost of equity = 0.1145 + 0.0167
Levered cost of equity = 0.1312 or 13.12%

WACC = Weight of debt * Cost of debt * (1 - Tax rate) + Weight of equity * Levered cost of equity
WACC = 0.28 * 7.15% * (1 - 0.21) + 0.72 * 13.12%
WACC = 1.58% + 9.45%
WACC = 11.03%